OpenStudy (anonymous):

Consider the following pair of equations: y = −2x + 8 y = x − 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

OpenStudy (lucaz):

substitute first y into the second equation -2x+8=x-1 and solve for x then take the value of x and subtitute into the first equation y=-2(x)+8

OpenStudy (anonymous):

Well, thank you so much. I really appreciate it!

OpenStudy (anonymous):

-2x+8=x-1 -2x-x = -1-8 -3x =-9 x = -9/-3 x = 3

OpenStudy (anonymous):

y = x − 1 y = 3 - 1 y = 2

OpenStudy (anonymous):

so (3,2)